Operations Report – January 3rd

SOL: 2
Name of person filing report: Mason Kuhn
Non-nominal systems: Hab toilet, EVA suit charger #1
Notes on non-nominal systems: No new issues with the toilet, just the sound that has been previously noted. EVA suit charger #1 has a voltage of 12.2 V whereas the rest are ~13.7 V. Suit 1 charge will not exceed 10.4 V, and 12.5 V is requested by mission support as the minimum for use on EVAs.
Spirit rover used: Yes
Hours: 206.4
Beginning charge: 100%
Ending charge: 67%
Currently charging: Yes
Opportunity rover used: Yes
Hours: 110.9
Beginning charge: 100%
Ending charge: 72%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 218.2 (educated estimate, forgot to record after training)
Beginning charge: 100%
Ending charge: 76% (educated estimate, forgot to record after training)
Currently charging: Yes
Perseverance rover used: Yes
Hours: 252.9 (educated estimate, forgot to record after training)
Beginning charge: 100%
Ending charge: 76%
Currently charging: Yes
General notes on rovers: Crew 272 had rover training this morning prior to entering Sim. All 4 rovers were used during training. EVA 1 from 1330 to 1445 used Spirit and Opportunity. EVA 2 from 1500 to 1615 used Spirit and Opportunity.
Summary of Hab operations: The damaged clothes hanging board in the 2nd stateroom from the right was fully repaired and is nominal.
WATER USE: 49.5 gallons
Water (static tank): 360.5 gallons
Static tank pipe heater (on or off): On
Static tank heater (On or off): On
Toilet tank emptied: Yes
Summary of internet: Nominal
Summary of suits and radios: Suits 2, 4, 5, 6, 7, 8, and 9 were used on EVAs. All suits are currently charging. All radios 1-7 were used during EVAs or as HabCom. They have been recharged and removed from the charging docks.
Summary of GreenHab operations:
WATER USE: 7 gallons
Heater: On
Supplemental light: 4 hours
Harvest: 20g beans, 3g basil, 3g chives, 6g dill, 28g carrots
Summary of ScienceDome operations:
Dual split: On
Summary of RAM operations: Retrieved a Phillips head screwdriver to repair a state room’s clothes hanging area. Replaced the screwdriver & DMM in the RAM at 1730. Took a measuring tape at 1730 to measure static tank water volume & put it back in the RAM immediately after.
Summary of any observatory issues: N/A
Summary of health and safety issues: One crew member got a small hand cut from an open can, and it was treated by washing, Neosporin, and a band-aid.
Questions, concerns, and requests to Mission Support:
Concern for Mission Support: Commander Kshitij is concerned that there is some airflow issue in the loft because his loft stateroom gets drafty and much colder than the rest of the Hab during the night.

Copyright © The Mars Society. All rights reserved. | Main Site